# Legal: Enhancing Legal Reasoning in LLMs via Reinforcement Learning with Chain-of-Thought Guided Information Gain

### v1

@@ Abstract
Legal Artificial Intelligence (LegalAI) has achieved notable advances in automating judicial decision-making with the support of Large Language Models (LLMs). However, existing legal LLMs still struggle to generate reliable and interpretable reasoning processes. They often default to fast-thinking behavior by producing direct answers without explicit multi-step reasoning, limiting their effectiveness in complex legal scenarios that demand rigorous justification. To address this challenge, we propose Legal, a reinforcement learning framework designed to enhance legal reasoning through chain-of-thought guided information gain. During training, Legal employs a dual-mode input setup—comprising direct answer and reasoning-augmented modes—and maximizes the information gain between them. This encourages the model to acquire meaningful reasoning patterns rather than generating superficial or redundant explanations. Legal follows a two-stage approach: (1) distilling latent reasoning capabilities from a powerful Large Reasoning Model (LRM), DeepSeek-R1, and (2) refining reasoning quality via differential comparisons, combined with a multidimensional reward mechanism that assesses both structural coherence and legal-domain specificity. Experimental results on multiple legal reasoning tasks demonstrate that Legal outperforms strong baselines in both accuracy and interpretability. It consistently produces more robust and trustworthy legal judgments without relying on labeled preference data. All code and data are available at https://github.com/NEUIR/LegalDelta.
@@ 1 Introduction
The rapid development of LegalAI systems—referring to artificial intelligence technologies applied to legal practice and judicial processes—is fundamentally reshaping the legal landscape. By automating complex tasks such as legal information extraction (Rabelo et al. 2022), intelligent case retrieval (Zeng et al. 2005; Zhang and Dou 2023), judgment prediction (Liu et al. 2023b), and legal question answering (Louis, van Dijck, and Spanakis 2024), LegalAI significantly enhances the efficiency and accessibility of legal services. These technologies have demonstrated substantial value in real-world applications, including contract analysis, legal dispute resolution, and regulatory compliance. Given these promising developments, recent research has increasingly turned to Large Language Models (LLMs) as a foundation for building a more effective legal assistant.
@@
Recently, researchers have increasingly adopted Reinforcement Learning (RL) frameworks such as Group Relative Policy Optimization (GRPO), which enables models to learn from feedback and optimize their behavior through trial-and-error. Large Reasoning Models (LRMs) like QwQ (Team 2025) and DeepSeek-R1 (Guo et al. 2025) have demonstrated strong reasoning capabilities across specialized domains including medicine (Sun et al. 2025), mathematics (Ahn et al. 2024), and programming (El-Kishky et al. 2025). Recent work has extended these ideas to the legal domain by distilling reasoning traces from powerful RL-trained models such as DeepSeek-V3 (Liu et al. 2024), as in LawGPT (Zhou et al. 2025). However, these methods typically rely on rule-based rewards in Reinforcement Learning with Verifiable Rewards (RLVR) (Mroueh 2025), which focus on output correctness while neglecting the quality and depth of the reasoning process itself. Therefore, how to effectively optimize Chain-of-Thought (COT) within the RL paradigm remains an open question.
@@
To address these challenges, this paper presents Legal, a reinforcement learning framework designed to improve legal reasoning in LLMs by incorporating Chain-of-Thought guided Information Gain. Built upon the GRPO paradigm, Legal introduces an information-gain based reward mechanism into the RLVR setting, encouraging the model to explore diverse reasoning trajectories and develop robust reasoning strategies. Specifically, Legal quantifies information gain by leveraging the mathematical equivalence between model logits and Q-values in reinforcement learning. By comparing model responses under direct inference and chain-of-thought prompting, it computes confidence differentials that capture both (i) pointwise mutual information between reasoning steps and final answer correctness, and (ii) global confidence shifts indicating whether reasoning improves decision reliability. This reward is then used to guide policy optimization, pushing the model to discover and reinforce high-quality reasoning behaviors that balance accuracy and prediction confidence.
Our experiments on various LLMs demonstrate that Legal outperforms all baseline models, achieving approximately 10% performance improvements across multiple legal reasoning tasks, while also exhibiting strong generalization to out-of-domain tasks. Further analysis reveals that the information gain module significantly enhances model confidence levels through progressive logits improvements, while also reducing training batch variance for improved training efficiency and stability. Additionally, comprehensive performance quality assessment demonstrates that Legal increases the prominence of legally tokens and generates reasoning chains with lower perplexity scores at the token-level, while consistently outperforming baseline models across different scenarios complexities at the sequence-level.
2 Related Work
Existing legal LLMs mainly rely on Supervised Fine-Tuning (SFT) to help models better understand legal semantics. These methods usually train the models on legal corpora using language modeling objectives. For example, models like LawGPT-zh (Liu et al. 2023a), HanFei (He et al. 2023), and Lawyer LLaMA (Huang et al. 2023) adopt SFT and combine it with the Low-Rank Adaptation (LoRA) strategy to improve training efficiency (Hu et al. 2022; Cui et al. 2023; Deng et al. 2023). One major challenge of SFT lies in the construction of high-quality training data. Recent work (Yue et al. 2023) addresses this by collecting existing publicly available legal datasets, crawling legal texts from the web, and incorporating instruction-tuning datasets (Zhang et al. 2023; Huang et al. 2023; Dai et al. 2023). While these approaches improve performance to some extent, they typically require extensive labeled data and focus on end-to-end optimization over task-specific labels. As a result, SFT-based methods are often susceptible to overfitting and catastrophic forgetting (Luo et al. 2023), which undermines the retention of previously acquired legal knowledge.
To address this challenge, recent studies (Verma et al. 2025; Vara et al. 2025) have explored Reinforcement Learning (RL) to encourage LLMs to engage in broader exploration by specially designed reward mechanisms. Intelligent Legal Assistant (Yao et al. 2025) utilizes the coverage of case elements as the reward signal to guide clarification in legal question answering. SyLeR (Zhang et al. 2025b) addresses the lack of structured reasoning paths in legal question answering by introducing a structure-aware similarity-based reward mechanism that guides LLMs to construct syllogistic reasoning paths. However, its reliance on surface-level matching, making it difficult to capture the complexity and rigor of legal reasoning. Simultaneously, Legal Mathematical Reasoning (Zhang et al. 2025a) employs result-oriented rewards to enhance mathematical reasoning, belonging to the Reinforcement Learning with Verifiable Rewards (RLVR) framework. However, this approach falls short in effectively evaluating intermediate reasoning quality. Therefore, we propose Legal that integrates an information-gain-enhanced reward mechanism into the RLVR framework to provide fine-grained supervision for reasoning quality, enabling LLM to learn more reliable and confidence-aligned legal reasoning strategies.
@@ 3 Methodology
In this section, we introduce the preliminaries of legal tasks (Section 3.1). Then we introduce the reinforcement learning based legal optimization method (Section 3.2). Finally, an information-gain-enhanced reward mechanism is proposed, which is based on differential Q-value analysis to assess reasoning quality (Section 3.3).
3.1 Preliminary of LLM Reasoning in the Legal Tasks
Legal reasoning is a complex cognitive process involving fact analysis, application of legal principles, and structured logical deduction. Unlike standard QA tasks, legal reasoning requires multi-step analytical thinking, integration of domain knowledge, and the generation of interpretable argumentation. Our framework focuses on five representative legal tasks that demand such capabilities.
Given a legal query , our model produces a structured output:
where denotes the step-by-step reasoning process and is the final answer. This structured output ensures both interpretability and reliability, which are crucial in legal applications. We then introduce our reinforcement learning based training strategy in Section 3.2, which equips LLMs with legal reasoning capabilities. Following this, Section 3.3 presents an information-theoretic reward enhancement mechanism that refines into via differential Q-value analysis, enabling more nuanced and reasoning-aware reward modeling for reinforcement learning.
@@ 3.2 Enhancing the Reasoning Capability of LLMs via Reinforcement Learning
General-domain LLMs often fall short in domain-specific reasoning tasks. Although supervised fine-tuning (SFT) (Gururangan et al. 2020) can adapt models to target domains, it also risks overfitting and catastrophic forgetting (Luo et al. 2023), thereby weakening general reasoning capabilities. To address this, we propose a RL-based training framework. The first stage employs SFT to guide the LLM to imitate the expert-like reasoning behaviors. The second stage applies reinforcement learning to further refine the reasoning ability based on performance feedback.
Supervised Fine-Tuning
We construct via SFT using reasoning-enhanced legal data. Specifically, we leverage the high-quality reasoning chains generated by DeepSeek-R1 (Guo et al. 2025) to construct a dataset pairing legal queries with structured and expert-like reasoning steps.
The SFT loss is defined as the standard cross-entropy loss over the full reasoning sequence:
where represents the training dataset, denotes the complete output containing both reasoning chain and final answer, and represents the probability of generating the complete reasoning output given the query and task instruction.
Reinforcement Learning
Building upon the distilled model, we apply Group Relative Policy Optimization (GRPO) (Shao et al. 2024) to further enhance the model’s reasoning ability. Unlike traditional RL approaches that require separate critic models, GRPO compares intra-group outputs to estimate advantages, making it effective for the scenarios where outputs can vary widely in quality.
@@
where , and is the group-wise advantage. The KL term penalizes policy shifts to maintain training stability.
The effectiveness of GRPO critically depends on the quality of reward signals. Legal defines a structured reward function as:
where ensures adherence to the structured output format, and evaluates the correctness of the legal reasoning process and conclusion.
Format Reward (): This component ensures outputs adhere to a structured format, with <reasoning> and <answer> tags separating the reasoning chain from the final decision. This format facilitates reward computation of both the reasoning process and the legal outcome.
Legal Reward (): This component evaluates the correctness and legality of the final answer using fine-grained, task-specific criteria. These legal criteria make the outcomes more verifiable and thus suitable for guiding GRPO-based training. Further details on the design of the legal reward are provided in Appendix A.3.
3.3 Information-Gain Enhanced Legal Reward Modeling
To move beyond terminal answer evaluation, we introduce an information-theoretic enhancement to reward design. Specifically, we replace the legal reward component with for reward modeling:
Recent work by Li et al. (Li et al. 2025) proves that logits from next-token prediction training are mathematically equivalent to Q-functions in offline inverse reinforcement learning(see detailed proof in Appendix A.4):
This formulation measures the semantic shift in the model’s internal representation space induced by the reasoning process , providing a quantitative measure of how much the reasoning contributes to the model’s confidence in generating the correct answer. We show our reasoning mode and direct mode prompt templates respectively in Appendix A.6.
where , is used to traverse all possible answer tokens. Through algebraic manipulation of the logit difference, we obtain:
This decomposition reveals the information-theoretic interpretation of our approach.
The first term represents the pointwise mutual information (PMI) between the answer token and the reasoning process , measuring the correlation between the reasoning chain and the specific answer: a positive value indicates that the reasoning process makes this answer more likely, a negative value means the reasoning actually decreases the probability of this answer, and zero implies no correlation.
The second term serves as a global confidence regularizer that captures whether the reasoning process enhances the model’s overall decisiveness: when positive, it indicates that has sharpened the probability distribution, making the model more confident overall; when negative, it suggests the reasoning has introduced confusion. Within the GRPO framework, since rewards are utilized for comparative evaluation across multiple generated reasoning trajectories, the global normalization constant remains invariant across all candidates for the same query, allowing us to focus on the information-theoretic signal that truly differentiates reasoning quality.
@@ 3.3 Information-Gain Enhanced Legal Reward Modeling
| Qwen2.5-3B-Instruct | ||||||||||||||
| Zero-shot | 62.37 | 65.52 | 50.10 | 44.85 | 55.10 | 34.80 | 55.80 | 52.65 | 47.46 | 46.55 | 74.71 | 70.94 | 74.18 | 62.77 |
| SFT (DiscLaw) | 68.09 | 64.25 | 43.24 | 38.38 | 45.50 | 39.20 | 46.00 | 49.24 | 38.98 | 43.20 | 61.76 | 53.75 | 57.82 | 51.10 |
| DPO | 64.93 | 65.02 | 50.42 | 43.33 | 59.50 | 52.40 | 56.40 | 56.00 | 45.76 | 46.55 | 74.12 | 71.88 | 73.45 | 62.35 |
| R1-Distill | 67.42 | 67.74 | 48.35 | 43.67 | 48.20 | 36.60 | 57.40 | 52.77 | 53.39 | 47.86 | 76.47 | 73.13 | 75.27 | 65.22 |
| Legal | 79.13 | 76.20 | 58.59 | 54.15 | 61.90 | 57.40 | 63.20 | 64.37 | 55.08 | 50.09 | 81.76 | 75.63 | 76.00 | 67.71 |
@@ 4 Experimental Methodology
4.1 Datasets
In our experiments, we select training data from portions of the CAIL2018 (Xiao et al. 2018), JEC_QA (Zhong et al. 2019), and LAIC2021 (Zhong et al. 2019) datasets, which encompass diverse legal reasoning scenarios across multiple domains. We first use 450 high-quality reasoning samples distilled from DeepSeek-R1 to establish foundational legal reasoning capabilities. Subsequently, for GRPO training of Legal, we collect 6,981 samples for the training set and 563 samples for the validation set across all tasks in our experiments. We select knowledge-intensive legal reasoning tasks for a comprehensive assessment, including legal article prediction (SPP-F), criminal charge prediction (CCP), sentencing prediction (SLP), similar case analysis (CAS) and financial amount calculation (CAC), which demand sophisticated reasoning over legal facts and regulations. To assess the generalization capability of our approach, we additionally evaluate performance on the out-of-domain DiscLaw benchmark, which contains various legal professional examinations distinct from our training domains. Detailed dataset statistics can be found in the Appendix A.1.
4.2 Baselines
In our experiments, we compare Legal with existing law models: Fuzi-7B (Deng et al. 2023), HanFei-7B (He et al. 2023), LexiLaw-6B (Li et al. 2024), ChatLaw-13B (Cui et al. 2023), and DiscLaw-13B (Yue et al. 2024). These law models vary in parameter scales and capabilities, all trained using various training methods on diverse legal datasets. To ensure fair and rigorous evaluation, we reimplemented a representative method (DiscLaw) based on the same Qwen2.5 series base models as our approach. Additionally, we compare with four baseline models: Zero-shot, R1-distill, DPO, and SFT (DiscLaw) for comprehensive evaluation. By evaluating LLMs across different scales, we can systematically assess the generalization ability of various training strategies.
4.3 Evaluation Metrics
We design comprehensive evaluation metrics for different task types following Zhou et al. (2025); Yue et al. (2023). For in-domain evaluation, we employ F1 and Jaccard (JAC) scores for SPP-F and CCP tasks, which assess the quality of multi-label classification. For SLP, CAS, and CAC tasks, we use accuracy scores. For out-of-domain evaluation, we assess generalization capability using accuracy scores across five tasks from the DiscLaw benchmark: PAE, NJE, PFE, UNGEE, and LBK.
4.4 Implementation Details
In our experiments, we adopt the Qwen2.5-Instruct series due to its superior reflective capabilities and cognitively informed reasoning behavior (Gandhi et al. 2025), including the 3B, 7B, and 14B parameter models (Qwen et al. 2025), as backbone language models to ensure a comprehensive evaluation across different model scales. During training, we use a batch size of 2 per GPU and a learning rate of 5e-5. For LoRA fine-tuning, we use a rank of 32. We use nucleus sampling with temperature of 0.9 and generate samples per prompt. For information-gain, we set a temperature of . More training implementation details of Legal are presented in Appendix A.2.
@@ 5 Evaluation Result
In this section, we first evaluate the performance of various legal LLMs on legal reasoning tasks, demonstrating Legal’s effectiveness across different model scales and its generalization to out-of-domain scenarios. We then analyze the effectiveness of our information-gain-enhanced reward modeling by examining logits variations during training and the correlation between information gain and reward values. Following this, we conduct a comprehensive performance quality assessment from both token-level and sequence-level perspectives: at the token level, we analyze legal token prominence progression and assess reasoning chain quality through perplexity evaluation; at the sequence level, we examine performance improvements across different scenarios complexities. Additionally, we provide some case studies in Appendix A.5.
@@ 5.1 Overall Performance
As shown in Table 1, Legal achieves consistently superior performance across different model scales and evaluation settings. On in-domain legal reasoning benchmarks, Legal outperforms all baselines, including Zero-shot, SFT (DiscLaw), DPO, and R1-Distill, with an average improvement of approximately 10% across Qwen2.5 models of varying capacities. While DPO benefits from preference optimization and R1-Distill leverages high-quality reasoning traces, Legal outperforms both, especially on complex reasoning tasks such as SLP and CAS.

### v3

@@ Abstract
Legal Artificial Intelligence (LegalAI) has achieved notable advances in automating judicial decision-making with the support of Large Language Models (LLMs). However, existing legal LLMs still struggle to generate reliable and interpretable reasoning processes. They often default to fast-thinking behavior by producing direct answers without explicit multi-step reasoning, limiting their effectiveness in complex legal scenarios that demand rigorous justification. To address this challenge, we propose Legal, a reinforcement learning framework designed to enhance legal reasoning through chain-of-thought guided information gain. During training, Legal employs a dual-mode input setup—comprising direct answer and reasoning-augmented modes—and maximizes the information gain between them. This encourages the model to acquire meaningful reasoning patterns rather than generating superficial or redundant explanations. Experimental results on multiple legal reasoning tasks demonstrate that Legal outperforms strong baselines and consistently produces more robust and trustworthy legal judgments without relying on labeled preference data. All code and data are available at https://github.com/NEUIR/LegalDelta.
Index Terms— Legal Reasoning, Reinforcement Learning
@@ 1 Introduction
The rapid development of LegalAI systems—referring to artificial intelligence technologies applied to legal practice and judicial processes—is fundamentally reshaping the legal landscape. By automating complex tasks such as legal information extraction [15], intelligent case retrieval [23], and judgment prediction [20], LegalAI has significantly improved the efficiency and accessibility of legal services, demonstrating substantial value in real-world applications such as contract analysis, legal dispute resolution, and regulatory compliance. Building on these advances, recent research has increasingly explored Large Language Models (LLMs) as a foundation for developing more effective legal assistants.
@@
Recently, researchers have increasingly adopted Reinforcement Learning (RL) frameworks such as Group Relative Policy Optimization (GRPO), which enables models to learn from feedback and optimize their behavior through trial-and-error. Large Reasoning Models (LRMs) like QwQ [18] and DeepSeek-R1 [6] have demonstrated strong reasoning capabilities across specialized domains including medicine [17], mathematics [1], and programming [4]. Recent work has extended these ideas to the legal domain by distilling reasoning traces from powerful RL-trained models such as DeepSeek-V3 [11], as in LawGPT. These methods typically rely on rule-based rewards in Reinforcement Learning with Verifiable Rewards (RLVR) [13], which primarily verify the correctness of outputs to optimize LRMs. However, the absence of more nuanced verifiable reward modeling limits their effectiveness in optimizing legal LLMs, as RLVR often overlooks the quality and depth of the reasoning process. Consequently, how to effectively optimize Chain-of-Thought (CoT) within the RL paradigm to achieve more convincing legal reasoning remains an open question.
@@
To address these challenges, this paper presents Legal, a reinforcement learning framework designed to improve legal reasoning in LLMs by incorporating Chain-of-Thought guided Information Gain. Built upon the GRPO paradigm, Legal introduces an information-gain based reward mechanism into the RLVR setting, encouraging LLMs to explore diverse reasoning trajectories and develop robust reasoning strategies. Specifically, Legal quantifies information gain by leveraging the mathematical equivalence between model logits and Q-values in reinforcement learning. By comparing model responses under direct inference mode and chain-of-thought prompting mode, it computes confidence differentials that capture both (i) pointwise mutual information between reasoning steps and final answer correctness, and (ii) global confidence shifts indicating whether reasoning improves decision reliability. This reward is then used to guide policy optimization, making Legal LLMs discover and reinforce high-quality reasoning behaviors that balance prediction accuracy and confidence.
Our experiments on various LLMs demonstrate the effectiveness of Legal, which outperforms all baseline models and achieves approximately 10% improvement across multiple legal reasoning tasks. Notably, Legal maintains its advantages on out-of-domain tasks, demonstrating strong generalization capability. Further analysis reveals that Legal effectively enhances model confidence through information gain-enhanced reward modeling, which progressively improves logits, increases the prominence of legal tokens, and further enhances reasoning quality.
@@ 2 Methodology
In this section, we introduce the reinforcement learning based legal optimization method (Section 2.1) and the information-gain-enhanced reward mechanism (Section 2.2).
@@ 2.1 Enhancing the Reasoning Capability of LLMs via Reinforcement Learning
Given a query , our model produces a structured output:
where denotes the step-by-step reasoning process and is the final answer. This structured output ensures both interpretability and reliability, which are crucial in legal applications.
General-domain LLMs often fall short in domain-specific reasoning tasks. Although Supervised Fine-Tuning (SFT) [7] can adapt models to target domains, it also risks overfitting and catastrophic forgetting [12], thereby weakening general reasoning capabilities. To address this, we propose a RL-based training framework. The first stage employs SFT to guide the LLM to imitate the expert-like reasoning behaviors. The second stage applies reinforcement learning to further optimize LLMs based on performance feedback.
Supervised Fine-Tuning. We construct via SFT using reasoning-enhanced legal data. Specifically, we leverage the high-quality reasoning chains generated by DeepSeek-R1 [6] to construct a dataset pairing legal queries with structured and expert-like reasoning steps. The SFT loss is defined as the standard cross-entropy loss over the full reasoning sequence:
where represents the training dataset and represents the probability of generating the complete output containing both reasoning chain and final answer given the query.
Reinforcement Learning. As shown in Figure 1, building upon the distilled model, we apply Group Relative Policy Optimization (GRPO) [16] to further enhance the model’s reasoning ability on these legal tasks. Unlike traditional RL approaches that require separate critic models, GRPO compares intra-group outputs to estimate advantages, making it effective for the scenarios where outputs can vary widely in quality.
@@
where , and is the group-wise advantage. The effectiveness of GRPO critically depends on the quality of reward signals. Legal defines a structured reward function as:
where the Format Reward () ensures outputs adhere to a structured format, with <reasoning> and <answer> tags separating the reasoning chain from the final answer. And the Legal Reward () evaluates the correctness and legality of the final answer using fine-grained, task-specific criteria.
@@
2.2 Information-Gain Enhanced Legal Reward Modeling
In this subsection, we introduce an information-theoretic enhancement to reward design. Specifically, we replace the legal reward component with in Eq. 4 for reward modeling:
Recent work by Li et al. [10] proves that logits from next-token prediction training are mathematically equivalent to Q-functions in offline inverse reinforcement learning:
This formulation measures the semantic shift in the model’s internal representation space induced by the reasoning process , providing a quantitative measure of how much the reasoning contributes to the model’s confidence in generating the correct answer.
where , is used to traverse all possible answer tokens. Then can be calculated:
This decomposition reveals the information-theoretic interpretation: The first term represents the pointwise mutual information (PMI) between the answer token and the reasoning process , measuring the correlation between the reasoning chain and the specific answer: a positive value indicates that the reasoning process makes this answer more likely, a negative value means the reasoning actually decreases the probability of this answer, and zero implies no correlation. The second term serves as a global confidence regularizer that captures whether the reasoning process enhances the model’s overall decisiveness: when positive, it indicates that has sharpened the probability distribution, making model more confident; else, it suggests the reasoning has introduced confusion.
@@ 3 Experimental Methodology
Datasets. We select training data from CAIL2018 [19] and JEC_QA [24] datasets. We first use 450 high-quality reasoning samples distilled from DeepSeek-R1 for foundational reasoning capabilities, then collect 6,981 training and 563 validation samples for GRPO training. We evaluate on five knowledge-intensive legal reasoning tasks: legal article prediction (SPP-F), criminal charge prediction (CCP), sentencing prediction (SLP), similar case analysis (CAS), and financial amount calculation (CAC). Additionally, we assess generalization on the DiscLaw benchmark [21].
Baselines. In our experiments, we compare Legal with existing law models: Fuzi-7B [3], HanFei-7B [8], LexiLaw-6B [9], ChatLaw-13B [2], and DiscLaw-13B [22]. These legal models vary in parameter scale and capability, and are trained using diverse methods on a range of legal datasets. To ensure fair and rigorous evaluation, we reimplemented a representative method SFT (DiscLaw) based on the same Qwen2.5 series base models in our experiments. Additionally, we compare with four baseline models: Zero-shot, R1-Distill, DPO, and SFT (DiscLaw) for a comprehensive evaluation.
Evaluation Metrics. Following recent work [21], we design comprehensive evaluation metrics: F1 and Jaccard (Jac) scores are used for SPP-F and CCP tasks, while accuracy is used for others.
Implementation Details. In our experiments, we adopt the Qwen2.5-Instruct series due to its superior reflective capabilities and cognitively informed reasoning behavior [5], including the 7B and 14B versions [14], as backbone LLMs to ensure a comprehensive evaluation across different model scales. During training, we use a batch size of 2 per GPU and a learning rate of 5e-5. For LoRA fine-tuning, we use a rank of 32. We use nucleus sampling with a temperature of 0.9 and generate samples per prompt. To calculate the information gain, we set a temperature of .
@@ 4 Evaluation Result
In this section, we evaluate the performance of various legal LLMs on legal reasoning tasks and then analyze the effectiveness of our information-gain-enhanced reward modeling.
@@ 4.1 Overall Performance
As shown in Table 1, Legal achieves consistently superior performance across different model scales and evaluation settings. On in-domain legal reasoning benchmarks, Legal outperforms all baselines, with an average improvement of approximately 10% across Qwen2.5 models of varying capacities. While DPO benefits from preference optimization and R1-Distill leverages high-quality reasoning traces, Legal outperforms both models, especially on complex reasoning tasks such as SLP and CAS.